OpenStudy (anonymous):

Write a linear factorization of the function. f(x) = x^4 + 36x^2 A. f(x) = x^2(6x + i)2 B. f(x) = x^2(x + 6i)2 C. f(x) = x^2(6x + i)(6x - i) D. f(x) = x^2(x + 6i)(x - 6i)

OpenStudy (anonymous):

Is it a?

OpenStudy (anonymous):

d

OpenStudy (anonymous):

What would you do with the 12i ?

OpenStudy (anonymous):

12i???

OpenStudy (anonymous):

Yea don't you get a 12i when u go backwards to get the same problem as before ?

OpenStudy (anonymous):

i will try to solve it please correct me

OpenStudy (anonymous):

x^4 +36x^2 = x^2(x^2 +6^2)

OpenStudy (anonymous):

and (x^2 +6^2) does not have solution on the reals numbers so

OpenStudy (anonymous):

x^2 +6^2= (x+6i)(x-6i)

OpenStudy (anonymous):

so the solution is x^2(x+6i)(x-6i)

OpenStudy (anonymous):

Ok I understand now I was looking at it in a different way :) thank u :)